OpenStudy (anonymous):

need help understanding how this was setup. int[0,1] pi(1-y^(2/3)dy. y= x^3????

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits_0^1 \pi(1-y^{2/3})\;dy\]\[\Large\bf\sf y=x^3\]I don't understand what the second line is. Is that supposed to be a substitution? Or is that the "solution"... or something?

OpenStudy (anonymous):

I figured it out. y^2/3 was the area after you subract it from the 1(the rest of the boxed in area. y=x^3 when evauated around the y-axis is x=y^1/3 thus r^2 = y^2/3...